HETAV PATEL

Grade 11,

The circles having radii r1 and r2 intersect orthogonally. Length of their common chord is

The circles having radii r1 and r2 intersect orthogonally. Length of their common chord is

Grade:12

3 Answers

Arun Kumar IIT Delhi
9 years ago
Two circles intersect orthogonally means r1 through point of contact makes right angle with r2 through that point


so the line intersecting has lenght
common chord length = h
by equating area of triangle formed in 2 different way

Arun Kumar
IIT Delhi
Askiitians Faculty
mayank rajput
7 years ago
you can solve this by using cosine formula in resulting triangle
if abc is a triangle then
C*2=A*2+B*2-2ABcosm
where m is angle between sides opposite to angle c.
 
Abhishek Mehta
7 years ago
you can solve this by using cosine formula in resulting triangleif abc is a triangle thenC*2=A*2+B*2-2ABcosmwhere m is angle between sides opposite to angle c.